# Fairness under Competition

### v1

@@ 3.1 Correlations between classifiers
Proof: For each $a$, the Pearson correlation between $B_{1}^{a}$ and $B_{2}^{a}$ is
@@
Proposition 1 yields the following simple corollary.
Corollary 1
We can also use Proposition 1 to analyze Example 1.
Example 1 continued
Suppose all classifiers in the example—the third-party classifier, as well as the two firms’ individual classifiers—have the same false-negative rate $\beta$. Thus, a firm that uses its own classifier on $A=1$ and the third-party classifier on $A=0$ is in practice using an EO classifier. Suppose also that the firms’ individual classifiers $c_{1}$ and $c_{2}$ are uncorrelated on $A=1$, namely, that $\rho^{1}=0$. Finally, note that, since both firms use the same classifier on $A=0$ we also have $\rho^{0}=1$. Thus, even though each firm uses an EO classifier, the level of EOC is $\sigma_{1}\sigma_{2}\left|{1-0}\right|=\beta(1-\beta)$.
3.2 Different sets of borrowers
Definition 1
$\Pr\left[c_{\ell}(X,A,Y)=0|Y=1,(X,A,Y)\in S_{1}\cap S_{2}\right]$
$~{}~{}~{}=\Pr\left[c_{\ell}(X,A,Y)=0|Y=1,(X,A,Y)\in S_{\ell}\setminus S_{3-\ell}\right],$
Proposition 2
To more clearly see the effect of different sets of served borrowers, consider the following corollary:
Corollary 2
Thus, a large EOC occurs when there is a large imbalance in the respective sizes of the overlaps in served borrowers between the two groups, $\gamma^{0}$ and $\gamma^{1}$. We now prove Proposition 2.
@@ 3.2 Different sets of borrowers
4 Harmful fairness adjustment
For the following examples, we assume that the fairness adjustment is derived via post-processing (Hardt et al., 2016): Given a learned classifier $c$, the EO classifier $\tilde{c}$ is derived from $c$, namely, it depends only on $A$ and the predictions $c(X,A)$. We also assume that $\tilde{c}$ minimizes squared-loss relative to all such derived EO classifiers. These simplifying assumptions imply the following lemma, a more formal version of which is stated and proved in the appendix:
Lemma 1
Example 3
$\Pr\left[c_{1}(X,A)=c_{2}(X,A)=0|Y=1,A=1\right]$
$\Pr\left[\tilde{c}_{1}(X,A)=\tilde{c}_{2}(X,A)=0|Y=1,A=1\right]$
$~{}~{}=\tilde{\rho}^{1}\cdot 0.1\cdot 0.9+0.1\cdot 0.1<0.1,$
where the inequality holds since $\tilde{\rho}^{1}<1$.
Example 4
5 Experiments
In order to explore the prevalence of our theoretical results we ran several experiments. For all experiments we used Lending Club loan data for the years 2007-2015, a dataset that includes roughly 890,000 peer-to-peer loans through Lending Club (2021). We used simple classifiers to predict the loan_status outcome, and, in particular, whether or not the loan was fully paid, given a set of features that included the loan amount, the interest rate, the number of installments, and the borrower’s annual income. As there is no demographic information in the public data, the protected attributes $A$ we used in our experiments were whether or not individuals have a mortgage. We compared the performance in terms of EO and EOC before and after the corresponding fairness adjustment. The fairness adjustment was implemented using the open-source Fairlearn (2025) package.
We ran four experiments. In the first three we simulated a situation in which two firms compete over the same pool of borrowers, $S_{1}=S_{2}$, as in Proposition 1. The differences between the firms’ classifiers were captured in three different ways in the respective experiments:
  • Exp. 1: Classifier $c_{1}$ was a logistic regression and $c_{2}$ a decision tree, but their training data was identical.
  • Exp. 2: Both classifiers were logistic regressions, but their training data was disjoint: one was trained on random examples with a short loan term, and one with a long loan term.
  • Exp. 3: Classifier $c_{1}$ was a logistic regression and $c_{2}$ a decision tree, and their training data was disjoint (as in Exp. 2).
In all three experiments, the training data consisted of a range of $300-100,000$ random examples. We ran each experiment with each size of training set 500 times. [^1] Running all experiments took a few hours on a home computer. For each run we calculated the EO of each classifier as well as the EOC, both before and after the fairness adjustments. The most relevant statistic was then the comparison between the EOC before fairness adjustments and the EOC after fairness adjustments. We calculated the percentage of times the EOC was higher after adjustment than before, as well as standard errors of the mean.
To examine the prevalence of this phenomenon we counted the percentage of times the EOC became worse after fairness adjustment. Figure 1 reports the results: The rows indicate the experiment number, the columns indicate the size of the training set, and the values in the table are the 95% confidence intervals, in percentages, for the probability that the EOC was worse after adjustment. We found this to be surprisingly common even for large training sets. For instance, for training sets of size 100k (out of a dataset of size roughly 900k) it occurred 30.0%, 15.6%, and 17.4% of the time in experiments 1, 2, and 3, respectively.
We note that the effect diminishes as the size of the training set increases. The diminishing of the effect for large training sets holds a similar message as the results from Section 3, where we show that the magnitude of the level of EOC is roughly the same as that of the false-negative rates. For very large training sets the false-negative rates diminish to 0, and so the levels of EO and EOC diminish in turn.
In the fourth experiment we simulated a situation in which $S_{1}\neq S_{2}$, as in Proposition 2. In particular, we simulated the situation described in Example 4, where $S_{1}=\mathrm{supp}(\mathcal{D})$ and $S_{2}=\mathrm{supp}(\mathcal{D}|A=1)$, and where $A=1$ is the set of individuals with a mortgage. Classifier $c_{1}$ was trained on a random set of examples taken from the entire dataset, whereas classifier $c_{2}$ was trained on a random set of examples consisting only of individuals with a mortgage. Both classifiers were logistic regressions, but since only classifier $c_{1}$ served individuals with both values of $A$, only that classifier underwent fairness adjustment. As in the previous experiments, we measured the level of EOC before and after the fairness adjustment.
Our results for experiment four are the following: For training sets of size up 10k, fairness adjustment leads to worse EOC in a significant fraction of runs. For larger training sets, however, the effect disappears. This is in line with the previous experiments, except that the effect diminishes for smaller training sets.
300 1k 3k 10k 30k 100k
6 Discussion and future work
In most of the paper we made simplifying assumptions on the number of lenders, the borrowers’ utility functions, and the fairness criterion used. In the technical appendix we extend some of our results to settings with more lenders and more general utility functions. Regarding the fairness criterion, our definition of WEC is general, but we leave the analysis of other special cases beyond EOC for future work. Our experimental results are also limited by our use of a single dataset and simple classifiers, and we leave further analysis of EOC in real-world settings to future work.
Appendix
Appendix A Extensions
In this section we consider two extensions of the simplified model presented in Section 2. First, in Section A.1, we study more general utility functions of the borrowers. And second, in Section A.2, we study competition between more than two classifiers.
A.1 Generalization: different utility functions
In this section we restrict again to two classifiers, but return to the more general form of the utility function, namely, $v:X\times A\times Y\times\mathbb{N}\mapsto{\mathbb{R}}$, where $\mathbb{N}$ is the set of non-negative integers. That is, $v(x,a,y,r)$ is the utility of a borrower $(x,a,y)$ when he receives $r$ loan offers. However, we now drop the assumption that $v(x,a,y,1)=v(x,a,y,2)$. Instead, we assume that $v(x,a,1,2)=k\cdot v(x,a,1,1)$ for some $k\geq 1$. As before, we assume that $v(x,a,y,0)=0$, and normalize $v(x,a,y,1)=1$. Our simplifying assumption in the main body of the paper was that $k=1$, but now we consider all values of $k\geq 1$. We first generalize Proposition 1 to this setting, and then apply it to Example 1.
Proposition 3
For two EO classifiers $c_{1}$ and $c_{2}$ with false-negative rates $\beta_{1}$ and $\beta_{2}$, the level of EOC is $\sigma_{1}\sigma_{2}\left|{(k-2)(\rho^{0}-\rho^{1})}\right|$. In the worst case, the level of EOC is $\left|{k-2}\right|\cdot\left(\min\{\beta_{1},\beta_{2}\}-\max\{0,\beta_{1}+\beta_{2}-1\}\right)$.
There are a few interesting observations to note. First, in all cases except $k=2$, the level of EOC is positive, even though both classifiers are EO. Hence, our result is robust. Interestingly, however, there is a qualitative difference between the case $k\in[1,2)$ and $k>2$, in that the group with lower utility under competition is different in the two parameter intervals. This last point is further explained at the end of this sub-section, in the context of the example.
Proof of Proposition 3 Recall from the proof of Proposition 1 that
and
Thus, we have that
$=1-\beta_{1}-\rho^{a}\sigma_{1}\sigma_{2}-(1-\beta_{1})(1-\beta_{2})$
$=\beta_{2}(1-\beta_{1})-\rho^{a}\sigma_{1}\sigma_{2}$
and, analogously,
Thus, for each $a$,
$\mathrm{E}\left[d(X,A)|Y=1,A=a\right]$ $=k\cdot\Pr\left[c_{1}(X,A)=c_{2}(X,A)=1|Y=1,A=a\right]$
$\quad+\Pr\left[c_{1}(X,A)\neq c_{2}(X,A)|Y=1,A=a\right]$
$=k\cdot\Pr\left[B_{1}^{a}=B_{2}^{a}=1\right]+\Pr\left[B_{1}^{a}\neq B_{2}^{a}\right]$
$=k\rho^{a}\sigma_{1}\sigma_{2}+k(1-\beta_{1})(1-\beta_{2})+\beta_{2}(1-\beta_{1})+\beta_{1}(1-\beta_{2})-2\rho^{a}\sigma_{1}\sigma_{2}$
$=(k-2)\rho^{a}\sigma_{1}\sigma_{2}+k-(k-1)(\beta_{1}+\beta_{2}-\beta_{1}\beta_{2}).$
The level of EOC is thus
$=\left|{(k-2)\rho^{0}\sigma_{1}\sigma_{2}-(k-2)\rho^{1}\sigma_{1}\sigma_{2}}\right|$
$=\sigma_{1}\sigma_{2}\left|{(k-2)(\rho^{0}-\rho^{1})}\right|.$
As in the proof of Proposition 1, the worst case scenario maximizes $\left|{\rho^{0}-\rho^{1}}\right|$, and so, for example, maximizes $\rho^{0}$ while minimizing $\rho^{1}$ (or vice versa). Again, as before, the worst-case difference is $\frac{1}{\sigma_{1}\sigma_{2}}\left(\min\{\beta_{1},\beta_{2}\}-\max\{0,\beta_{1}+\beta_{2}-1\}\right)$, which leads to the claimed worst case level of EOC.
Proposition 3 yields the following simple corollary.
Corollary 3
If $\beta=\beta_{1}=\beta_{2}$ and $\beta\leq 1/2$, then the worst-case level of EOC is $\beta\cdot\left|{k-2}\right|$.
We can also use Proposition 3 to further analyze Example 1.
Example 1 continued
Suppose again that all classifiers in the example—the third-party classifier, as well as the two firms’ individual classifiers—have the same false-negative rate $\beta$. Thus, a firm that uses its own classifier on $A=1$ and the third-party classifier on $A=0$ is in practice using an EO classifier. Suppose also that the firms’ individual classifiers $c_{1}$ and $c_{2}$ are uncorrelated on $A=1$, namely, that $\rho^{1}=0$. Finally, note that, since both firms use the same classifier on $A=0$ we also have $\rho^{0}=1$. Thus, even though each firm uses an EO classifier, the level of EOC is $\beta(1-\beta)\cdot{k-2}$. Furthermore, the utility of borrowers in group $A=0$ is $k(1-\beta)$, since all borrowers who receive an offer actually receive two offers. In group $A=1$, on the other hand, the expected utility of borrowers is $k(1-\beta)^{2}+2\beta(1-\beta)$.
Now,
This implies that the level of EOC is $\beta(1-\beta)\left|{2-k}\right|$. However, note that, if $k\in[1,2)$ then the difference is positive, whereas if $k>2$ then the difference is negative. This implies that, in the former case, the expected utility is higher in group $A=1$, whereas in the latter case, the expected utility is higher in group $A=0$.
The main insights from this section are, first, that two EO classifiers can lead to a positive EOC even for more general utility functions, and second, that whether or not the “disadvantaged” group is the one that suffers under a positive EOC depends on the utility function (i.e., whether $k\in[1,2)$ or $k>2$).
A.2 Generalization: more than two classifiers
Suppose that instead of two classifiers, the set $L$ contains $n$ classifiers. We provide two results. First, we characterize the worst-case EOC with $n$ classifiers. Second, we generalize the analysis of Example 1 from Section 3 and show that, when classifiers are uncorrelated on one group but fully correlated on the other group, the EOC strictly increases with $n$.
Suppose there are $n$ classifiers, $c_{1},\ldots,c_{n}$,
Proposition 4
Proof: The worst case occurs when the probability that all classifiers misclassify positive instances is minimal in group $A=1$ and maximal in group $A=0$ (or vice versa). To maximize $pr{c_{1}(X,A)=\ldots=c_{n}(X,A)=0|A=0,Y=1}$ the overlap in instances on which each $c_{i}$ misclassifies has to be maximal, which occurs when the instances are contained in one anther. Formally, if $\beta_{1}\leq\ldots\leq\beta_{n}$, then
$c_{n}(x,0)=1\Rightarrow\ldots\Rightarrow c_{1}(x,0)=1.$
In this case,
$pr{c_{1}(X,A)=\ldots=c_{n}(X,A)=0|A=0,Y=1}=\beta_{n}=\min_{i\in L}\beta_{i}.$
To minimize $pr{c_{1}(X,A)=\ldots=c_{n}(X,A)=0|A=1,Y=1}$ the overlap in instances on which each $c_{i}$ misclassifies has to be minimal, which occurs when the instances are maximally disjoint. Equivalently, the overlap in positive instances in which the $c_{i}$’s correctly classify is maximally disjoint. If $\sum_{j}(1-\beta_{j})\geq 1$ then
$\{(x,1,1):c_{1}(x,1)=\ldots=c_{n}(x,1)=0\}=\emptyset.$
Example 1 continued
$\Pr\left[c_{1}(X,A)=\ldots=c_{n}(X,A)=0|A=1,Y=1\right]=\beta^{n}.$
Appendix B Proof of Lemma 1
Recall the assumption that the fairness adjustment is derived via post-processing (Hardt et al., 2016): Given a learned classifier $c$, the EO classifier $\tilde{c}$ is derived from $c$, namely, it depends only on $A$ and the predictions $c(X,A)$. We also assume that $\tilde{c}$ minimizes squared-loss relative to all such derived EO classifiers. Under these simplifying assumptions we prove the following lemma:
Lemma 2
Proof: For a classifier $c$ and a group $a\in A$, let $\lambda^{a}(c)=\left(\alpha^{a},1-\beta^{a}\right)$, where $\alpha^{a}=\Pr\left[c_{L}(X,A,Y)=1|Y=0,A=a\right]$ is the false-positive rate of the classifier in group $a$, and $1-\beta^{a}=\Pr\left[c_{L}(X,A,Y)=1|Y=1,A=a\right]$ is the true-positive rate of the classifier in group $a$. Hardt et al. (2016) show that any derived classifier $\tilde{c}$ satisfies $\lambda^{a}(\tilde{c})\in\mathrm{convhull}\left\{(0,0),\lambda^{a}(c),\lambda^{a}(1-c),(1,1)\right\}$, where $(1-c)$ is the classifier $c$ but with predictions flipped. They then show that the optimal derived classifier can be found using the following linear program:

### v2

@@ 3.1 Correlations between classifiers
The intuition is the following. Under high (resp., low) correlation, if a borrower is misclassified by one classifier, she is likely also (resp., not) misclassified by the other. In the low correlation case, each deserving borrower intuitively has “two chances” to get an offer, whereas in the high correlation case she only has “one chance”. If in one group deserving borrowers get two chances and in the other they only get one, then in the former there is a lower chance of being misclassified.
The proofs of this and all other propositions are deferred to Appendix A. Proposition 1 yields the following simple corollary.
Corollary 1
We can also use Proposition 1 to analyze Example 1.
Example 1 continued
Suppose all classifiers in the example—the third-party classifier, as well as the two firms’ individual classifiers—have the same false-negative rate $\beta$. Thus, a firm that uses its own classifier on $A=1$ and the third-party classifier on $A=0$ is in practice using an EO classifier. Suppose also that the firms’ individual classifiers $c_{1}$ and $c_{2}$ are uncorrelated on $A=1$, namely, that $\rho^{1}=0$. Finally, note that, since both firms use the same classifier on $A=0$ we also have $\rho^{0}=1$. Thus, even though each firm uses an EO classifier, the level of EOC is $\sigma_{1}\sigma_{2}\left|{0-1}\right|=\beta(1-\beta)$.
3.1.1 Generalization: $v$-EOC
Recall that, under Assumption 1, EOC is equivalent to $v$-EOC, in which case all our results apply to the latter notion as well. In this section we study $v$-EOC when utility functions do not satisfy Assumption 1. In particular, we drop the assumption that $v(x,a,y,1)=v(x,a,y,2)=1$, and instead assume the following:
Assumption 2 (0-1-k preferences)
$v(x,a,y,r)=k\geq 1$ for $r>1$, $v(x,a,y,1)=1$, and $v(x,a,y,0)=0$, $\forall x,a,y$.
Assumption 1 imposes $k=1$, but now we consider all values of $k\geq 1$. Observe that under Assumption 2, a classifier is still EO if and only if it is $v$-EO. However, for a pair of classifiers $c_{1}$ and $c_{2}$, the definitions of EOC and $v$-EOC are no longer equivalent.
We first generalize Proposition 1 to this setting, and then apply it to Example 1.
Proposition 2
Under Assumption 2, for two EO classifiers $c_{1}$ and $c_{2}$ with false-negative rates $\beta_{1}$ and $\beta_{2}$, the level of $v$-EOC is $\sigma_{1}\sigma_{2}\left|{(k-2)(\rho^{0}-\rho^{1})}\right|$. In the worst case, the level of $v$-EOC is $\left|{k-2}\right|\cdot\left(\min\{\beta_{1},\beta_{2}\}-\max\{0,\beta_{1}+\beta_{2}-1\}\right)$.
There are a few interesting observations to note. First, in all cases except $k=2$, the level of $v$-EOC is positive, even though both classifiers are EO (and so, under Assumption 2, also $v$-EO). Hence, our insight is, in a sense, robust. Interestingly, however, there is a qualitative difference between the case $k\in[1,2)$ and $k>2$, in that the group with lower utility under competition is different in the two parameter intervals. This last point is further explained at the end of this sub-section, in the context of the example.
Proposition 2 yields the following simple corollary.
Corollary 2
If $\beta=\beta_{1}=\beta_{2}$ and $\beta\leq 1/2$, then the worst-case level of $v$-EOC is $\beta\cdot\left|{k-2}\right|$.
We can also use Proposition 2 to further analyze Example 1.
Example 1 continued
Suppose again that all classifiers in the example—the third-party classifier, as well as the two firms’ individual classifiers—have the same false-negative rate $\beta$. Thus, a firm that uses its own classifier on $A=1$ and the third-party classifier on $A=0$ is in practice using an EO classifier. Suppose also that the firms’ individual classifiers $c_{1}$ and $c_{2}$ are uncorrelated on $A=1$, namely, that $\rho^{1}=0$. Finally, note that, since both firms use the same classifier on $A=0$ we also have $\rho^{0}=1$. Thus, even though each firm uses an EO classifier, the level of $v$-EOC is $\beta(1-\beta)\cdot{k-2}$. Furthermore, the utility of borrowers in group $A=0$ is $k(1-\beta)$, since all borrowers who receive an offer actually receive two offers. In group $A=1$, on the other hand, the expected utility of borrowers is $k(1-\beta)^{2}+2\beta(1-\beta)$.
Now,
This implies that the level of $v$-EOC is $\beta(1-\beta)\left|{2-k}\right|$. However, note that, if $k\in[1,2)$ then the difference is positive, whereas if $k>2$ then the difference is negative. This implies that, in the former case, the expected utility is higher in group $A=1$, whereas in the latter case, the expected utility is higher in group $A=0$.
The main insights from this section are, first, that two $v$-EO classifiers can lead to a positive $v$-EOC even for more general utility functions, and second, that whether or not the “disadvantaged” group is the one that suffers under a positive $v$-EOC depends on the utility function (i.e., whether $k\in[1,2)$ or $k>2$).
3.1.2 Generalization: more than two classifiers
Suppose that instead of two classifiers, the set $L$ contains $n$ classifiers. We provide two results. First, we characterize the worst-case EOC with $n$ classifiers. Second, we generalize the analysis of Example 1 and show that, when classifiers are uncorrelated on one group but fully correlated on the other group, the level of EOC strictly increases (and so worsens) with $n$.
Proposition 3
Example 1 continued
$\Pr\left[c_{1}(X,A)=\ldots=c_{n}(X,A)=0|A=1,Y=1\right]=\beta^{n}.$
3.2 Different sets of borrowers
Definition 1
$\Pr\left[c_{\ell}(X,A)=0|Y=1,(X,A,Y)\in S_{1}\cap S_{2}\right]$
$~~~=\Pr\left[c_{\ell}(X,A)=0|Y=1,(X,A,Y)\in S_{\ell}\setminus S_{3-\ell}\right],$
Proposition 4
The intuition for Proposition 4 is similar to that of Proposition 1, with high (resp., low) overlap in the former playing the same role as high (resp., low) correlation in the latter. To more clearly see the effect of different sets of served borrowers, consider the following corollary:
Corollary 3
Thus, a large EOC occurs when there is a large imbalance in the respective sizes of the overlaps in served borrowers between the two groups, $\gamma^{0}$ and $\gamma^{1}$.
4 Harmful fairness adjustment
For the following examples, we assume that the fairness adjustment is derived via post-processing (Hardt et al., 2016): Given a learned classifier $c$, the EO classifier $\tilde{c}$ is derived from $c$, namely, it depends only on $A$ and the predictions $c(X,A)$. We also assume that $\tilde{c}$ minimizes squared-loss relative to all such derived EO classifiers. These simplifying assumptions imply the following lemma:
Lemma 1
Example 3
$\Pr\left[c_{1}(X,A)=c_{2}(X,A)=0|Y=1,A=1\right]$
$\Pr\left[\tilde{c}_{1}(X,A)=\tilde{c}_{2}(X,A)=0|Y=1,A=1\right]$
$~~=\tilde{\rho}^{1}\cdot 0.1\cdot 0.9+0.1\cdot 0.1<0.1,$
where the inequality holds since $\tilde{\rho}^{1}<1$.
Example 4
5 Experiments
In order to explore the prevalence of our theoretical results we ran several experiments. [^3] Code is available at https://github.com/eilamshapira/FairnessUnderCompetition. We describe and report results from the first three here, and the remaining ones in Appendix C. For the first three we used Lending Club loan data for the years 2007-2015, a dataset that includes roughly 890,000 peer-to-peer loans through Lending Club (2021). We used simple classifiers to predict the loan_status outcome, and, in particular, whether or not the loan was fully paid, given a set of features that included the loan amount, the interest rate, the number of installments, and the borrower’s annual income. As there is no demographic information in the public data, the protected attributes $A$ we used in our experiments were whether or not individuals have a mortgage. We compared the performance in terms of EO and EOC before and after the corresponding fairness adjustment. The fairness adjustment was implemented using the open-source Fairlearn (2025) package.
In the experiments we simulated a situation in which two firms compete over the same pool of borrowers, $S_{1}=S_{2}$, as in Proposition 1. The differences between the firms’ classifiers were captured in three different ways in the respective experiments:
  • Exp. 1
    $c_{1}$ was a logistic regression and $c_{2}$ a decision tree, but their training data was identical.
  • Exp. 2
    Both classifiers were logistic regressions, but their training data was disjoint: one was trained on random examples with a short loan term, and one with a long loan term.
  • Exp. 3
    $c_{1}$ was a logistic regression and $c_{2}$ a decision tree, and their training data was disjoint (as in Exp. 2).
In all three experiments, the training data consisted of a range of 300100,000300-100,000 random examples. We ran each experiment with each size of training set 500 times. [^4] Running all experiments took a few hours on a home computer. For each run we calculated the EO of each classifier as well as the EOC level, both before and after the fairness adjustments. The most relevant statistic was then the comparison between the EOC level before fairness adjustments and the EOC level after fairness adjustments. We calculated the percentage of times the EOC level was higher after adjustment than before, as well as standard errors of the mean.
To examine the prevalence of this phenomenon we counted the percentage of times the EOC became worse after fairness adjustment. Table 1 reports the results: The rows indicate the experiment number, the columns indicate the size of the training set, and the values in the table are the 95% confidence intervals, in percentages, for the probability that the EOC level was worse after adjustment. We found this to be surprisingly common even for large training sets. For instance, for training sets of size 100k (out of a dataset of size roughly 900k) it occurred 30.1%, 15.8%, and 17.4% of the time in experiments 1, 2, and 3, respectively.
We note that the effect diminishes as the size of the training set increases. The diminishing of the effect for large training sets holds a similar message as the results from Section 3, where we show that the magnitude of the EOC level is roughly the same as that of the false-negative rates. For very large training sets the false-negative rates diminish to 0, and so the levels of EO and EOC diminish in turn.
In addition to examining the prevalence of the effect, we also measured its magnitude. For each run, if the EOC level was higher after fairness adjustment than before, we measured the factor by which the EOC level increased. For each experiment we then calculated the average factor by which the EOC level increased (conditional on increasing), as well as the standard error of the mean. The results are reported in Appendix C. However, to get a sense of the orders-of-magnitude consider the following. For training sets of size 100k (out of a dataset of size roughly 900k), the EOC level increased by a mean factor of 19, 1.3, and 3.1 in experiments 1, 2, and 3, respectively.
In Appendix C we also report on additional experiments we ran. In particular, we ran an experiment with $S_{1}\neq S_{2}$, as in Proposition 4, and simulated the situation described in Example 4. We also ran experiments in which classifiers’ training sets were independently chosen, experiments using a different dataset (specifically, Becker and Kohavi, 1996), and experiments with three rather than two competing classifiers. For all experiments, the results were similar to the ones reported here.
300 1k 3k 10k 30k 100k
6 Discussion and future work
Our main insight is that individual fairness is neither necessary nor sufficient for ecosystem fairness. Thus, when making fairness adjustments, their effects on the ecosystem should be considered. An interesting question relates to the kinds of regulations that could lead to ecosystem fairness while also maintaining the benefits of having multiple competing classifiers. [^5] Dividing responsibility, for example by allowing only one EO classifier to serve each borrower, would lead to ecosystem fairness, but would not deliver the benefits of having multiple classifiers.
In our analysis we made several simplifying assumptions. First, we focused on EO and EOC. In Appendix B, however, we extend the definitions and analyses to two other common notions of fairness—namely, Equalized Odds and Demographic Parity. In particular, we define variants of these notions for a setting with multiple classifiers, and then show that versions of Propositions 1 and 4 hold for them as well: If the correlations or overlaps in served borrowers between classifiers differ across groups, then classifiers that are fair will not be fair under competition.
Second, for our welfare-based notions of $v$-EO and $v$-EOC, we focused on simple utility functions that satisfy Assumptions 1 or 2. A natural direction that we leave for future work is to consider more general preferences, including ones where deserving borrowers’ utilities may differ from others’, and where the utility function may depend on any of $(x,a,y)$.
Acknowledgments and Disclosure of Funding
Research supported by funding from the Israel Science Foundation, award number 2039/24 (Gradwohl), and by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant number 740435 (Tennenholtz).
Appendix
Appendix A Proofs
Proof of Proposition 1 For each $a$, the Pearson correlation between $B_{1}^{a}$ and $B_{2}^{a}$ is
@@ Appendix A Proofs
Proof of Proposition 2 Recall from the proof of Proposition 1 that
and
Thus, we have that
$=1-\beta_{1}-\rho^{a}\sigma_{1}\sigma_{2}-(1-\beta_{1})(1-\beta_{2})$
$=\beta_{2}(1-\beta_{1})-\rho^{a}\sigma_{1}\sigma_{2}$
and, analogously,
Thus, for each $a$,
$\mathrm{E}\left[v(X,A,Y,R(X,A))|Y=1,A=a\right]=k\cdot\Pr\left[c_{1}(X,A)=c_{2}(X,A)=1|Y=1,A=a\right]$
$\quad\quad+\Pr\left[c_{1}(X,A)\neq c_{2}(X,A)|Y=1,A=a\right]$
$\quad=k\cdot\Pr\left[B_{1}^{a}=B_{2}^{a}=1\right]+\Pr\left[B_{1}^{a}\neq B_{2}^{a}\right]$
$\quad=k\rho^{a}\sigma_{1}\sigma_{2}+k(1-\beta_{1})(1-\beta_{2})+\beta_{2}(1-\beta_{1})+\beta_{1}(1-\beta_{2})-2\rho^{a}\sigma_{1}\sigma_{2}$
$\quad=(k-2)\rho^{a}\sigma_{1}\sigma_{2}+k-(k-1)(\beta_{1}+\beta_{2}-\beta_{1}\beta_{2}).$
The level of $v$-EOC is thus
$\left|{\mathrm{E}\left[v(X,A,Y,R(X,A))|Y=1,A=0\right]-\mathrm{E}\left[v(X,A,Y,R(X,A))|Y=1,A=1\right]}\right|$
$\quad=\left|{(k-2)\rho^{0}\sigma_{1}\sigma_{2}-(k-2)\rho^{1}\sigma_{1}\sigma_{2}}\right|$
$\quad=\sigma_{1}\sigma_{2}\left|{(k-2)(\rho^{0}-\rho^{1})}\right|.$
As in the proof of Proposition 1, the worst case scenario maximizes $\left|{\rho^{0}-\rho^{1}}\right|$, and so, for example, maximizes $\rho^{0}$ while minimizing $\rho^{1}$ (or vice versa). Again, as before, the worst-case difference is $\frac{1}{\sigma_{1}\sigma_{2}}\left(\min\{\beta_{1},\beta_{2}\}-\max\{0,\beta_{1}+\beta_{2}-1\}\right)$, which leads to the claimed worst case level of $v$-EOC.
Proof of Proposition 3 The worst case occurs when the probability that all classifiers misclassify positive instances is minimal in group $A=1$ and maximal in group $A=0$ (or vice versa). To maximize $\Pr\left[c_{1}(X,A)=\ldots=c_{n}(X,A)=0|A=0,Y=1\right]$ the overlap in instances on which each $c_{i}$ misclassifies has to be maximal, which occurs when the instances are contained in one anther. Formally, if $\beta_{1}\leq\ldots\leq\beta_{n}$, then
$c_{n}(x,0)=1\Rightarrow\ldots\Rightarrow c_{1}(x,0)=1.$
In this case,
$\Pr\left[c_{1}(X,A)=\ldots=c_{n}(X,A)=0|A=0,Y=1\right]=\beta_{n}=\min_{i\in L}\beta_{i}.$
To minimize $\Pr\left[c_{1}(X,A)=\ldots=c_{n}(X,A)=0|A=1,Y=1\right]$ the overlap in instances on which each $c_{i}$ misclassifies has to be minimal, which occurs when the instances are maximally disjoint. Equivalently, the overlap in positive instances in which the $c_{i}$’s correctly classify is maximally disjoint. If $\sum_{j}(1-\beta_{j})\geq 1$ then
$\{(x,1,1):c_{1}(x,1)=\ldots=c_{n}(x,1)=0\}=\emptyset.$
@@
Proof of Lemma 1 For a classifier $c$ and a group $a\in A$, let $\lambda^{a}(c)=\left(\alpha^{a},1-\beta^{a}\right)$, where $\alpha^{a}=\Pr\left[c_{L}(X,A,Y)=1|Y=0,A=a\right]$ is the false-positive rate of the classifier in group $a$, and $1-\beta^{a}=\Pr\left[c_{L}(X,A,Y)=1|Y=1,A=a\right]$ is the true-positive rate of the classifier in group $a$. Hardt et al. (2016) show that any derived classifier $\tilde{c}$ satisfies $\lambda^{a}(\tilde{c})\in\mathrm{convhull}\left\{(0,0),\lambda^{a}(c),\lambda^{a}(1-c),(1,1)\right\}$, where (1c)(1-c) is the classifier $c$ but with predictions flipped. They then show that the optimal derived classifier can be found using the following linear program:
@@
$\textup{s.t.}\quad\lambda^{a}(\tilde{c})\in\mathrm{convhull}\left\{(0,0),\lambda^{a}(c),\lambda^{a}(1-c),(1,1)\right\},\quad\forall a\in A$